a student is going to school. she makes it to school with an average speed of 35 miles per hour. the drive home is the exact same distance, but the average speed is 40 miles per hour. what is the average speed of the round trip?
Answer by Alan3354(69443)

a student is going to school. she makes it to school with an average speed of 35 miles per hour. the drive home is the exact same distance, but the average speed is 40 miles per hour. what is the average speed of the round trip?
-------
Avg = 2*35*40/(35+40) = 2800/75 = 112/3
= 37.33... mi/hr